(ii) Given the following Cobb Douglas Utility function u(x1, x2) = x1cx2d.What is Marginal Rate of Substitution MRS x1, x2 ?



1
Expert's answer
2021-07-01T13:18:45-0400

Need to find-

Marginal Rate of Substitution "MRS x_1, x_2"

Given in the question-

"u(x_1, x_2) = x_1^{c}x_2^{d}"

"MRS =\frac{ MU_x1}{ MU_x2}................(1)\\

MU_x1 =\frac{\delta u(x_1, x_2)}{\delta x_1}\\

MU_x1 = \frac{\delta(x_1^{c}x_2^{d})}{\delta x_1}\\

MU_x1 = c\times(x1^{c-1})\times(x2^d )"

similarly,

"MU_x2 = \delta(x1^c)\times(x2^{d-1})"

So, from eqn."1"

"MRS = (c\times(x1^{c-1})\times(x2^{d} ))\div(\delta(x1^c)\times(x2^{d-1}))"

After simplifying 

"MRS = c(x2)\div d(x1)"